# An Elliott intertwining approach to classifying actions of $\mathrm{C}^{*}$-tensor categories

### 29 Jan 2024 v2

@@ Acknowledgements
We would like to thank Stuart White and Samuel Evington for their supervision on this project. We would also like to thank George Elliott and Gabór Szabó for useful discussions on the topic of this paper. Part of this work was completed during the authors’ stay at the Fields Institute for Research in Mathematical Sciences for the ’Thematic Program on Operator Algebras and Applications’ in Autumn 2023. We thank the Fields Institute and the organisers for the hospitality.
The first named author was supported by the Ioan and Rosemary James Scholarship awarded by St John’s College and the Mathematical Institute, University of Oxford, as well as by project G085020N funded by the Research Foundation Flanders (FWO). The second named author was supported by the EPSRC grant EP/R513295/1. The authors’ stay at the Fields Institute was partially funded by a Special Grant awarded by St John’s College, Oxford. The first named author’s stay was also partially supported by the Fields Institute while the travel costs of the second named author were supported by the EPSRC grant EP/X026647/1.

### 08 Apr 2026 v3

@@ Acknowledgements
We would like to thank Stuart White and Samuel Evington for their supervision on this project. We would also like to thank Corey Jones, George Elliott, and Gabór Szabó for useful discussions on the topic of this paper. Part of this work was completed during the authors’ stay at the Fields Institute for Research in Mathematical Sciences for the ’Thematic Program on Operator Algebras and Applications’ in Autumn 2023. We thank the Fields Institute and the organisers for the hospitality.
The first named author was supported by the Ioan and Rosemary James Scholarship awarded by St John’s College and the Mathematical Institute, University of Oxford, as well as by projects G085020N and 1249225N funded by the Research Foundation Flanders (FWO). The second named author was supported by the EPSRC grant EP/R513295/1, by the European Research Council under the European Union’s Horizon Europe research and innovation programme (ERC grant AMEN–101124789), and by the postdoctoral fellowship 1204626N of the Research Foundation Flanders (FWO).
The authors’ stay at the Fields Institute was partially funded by a Special Grant awarded by St John’s College, Oxford. The first named author’s stay was also partially supported by the Fields Institute while the travel costs of the second named author were supported by the EPSRC grant EP/X026647/1.
For the purpose of open access, the authors have applied a $C$ BY public copyright license to any author accepted manuscript version arising from this submission.